OpenStudy (anonymous):

What is the limit of (cos pi x)/3 as x approaches 2?

OpenStudy (mertsj):

\[\lim_{x \rightarrow 2}\frac{\cos \pi x}{3}=\frac{\cos 2\pi}{3}=\frac{1}{3}\]